OpenStudy (anonymous):

Hello can you please help me? I understand if you can't. Question: is 4 part of the solution set for the inequality 7w < (line under the <) 28? why or why not?

OpenStudy (anonymous):

\[7w\leq 28\] \[w\leq 4\] by dividing by 7

OpenStudy (anonymous):

so yes, 4 is part of the solution set because \[4\leq 4\] is a true statement

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

yw

OpenStudy (radar):

@satellite73 what is with the "Math Processing Error" in red associated with your post?

OpenStudy (anonymous):

how can i say > (line under >)

OpenStudy (anonymous):

oh lord not again

OpenStudy (anonymous):

@radar with this new mathjax sometimes you cannot see it. apparently xSOS can though. maybe try refreshing, it was out for me for a while

OpenStudy (radar):

"greater than or equal to" just GTE will convey the meaning.

OpenStudy (anonymous):

@xSOScandiiex you read \[\leq\] as "less than or equal to" how you type it is \leq